# Best Practices for Multi-Fidelity Bayesian Optimization in Materials and Molecular Research

### v1

@@ Abstract
Although MFBO has garnered interest in the past decade within the ML community, leading to a myriad of model definitions and acquisition policies 9, 10, 11, researchers have only lately started to integrate them within their design methodologies. In the chemical domain, there has been recent interest in incorporating cost awareness into the BO loop12. In the specific multi-fidelity approach, several studies have successfully applied MFBO to the materials discovery domain 13, 14, 15, 16. However, there is a fundamental lack in the assessment of MFBO performance, and due to this, each work represents MFBO performance in different ways 9, 11, 14. Although some metrics have been proposed 13, the lack of clarity and unified criteria on how to assess the benefits of MFBO in the chemical domain has hindered its widespread adoption among the experimental community17. Several factors come into play when assessing its impact, yet the application of MFBO can also be detrimental to the overall optimization process 18. It has been also shown how in the long run the advantage of MFBO over SFBO can be lost 17.
@@
Figure 1: Multi-fidelity Bayesian Optimization (MFBO) combines expensive but informative and cheap but approximate sources of information with a ML model to optimize black-box problems at an overall reduced budget. The high-fidelity source (HF, e.g., real lab experiment) provides accurate information but at a high cost. The low-fidelity source (LF, e.g., computer simulation of real experiment) represents a cheaper approximation of the HF. In the illustration, we show the surfaces corresponding to the Branin function used in this study at the HF and one LF level, where the optima are located in different places. The MFBO loop iteratively selects HF or LF to maximize the information gain while reducing the overall optimization cost.
@@ Methods
where f is the target black-box function and $\mathcal{X}$ is the space of all possible input candidates. Considering $\mathcal{D}=[(x_{1},y_{1}),(x_{2},y_{2}),...(x_{n},y_{n})]$ a dataset containing pairs of inputs-outputs $({x}_{n},y_{n})\in\mathbb{R}^{d}\times\mathbb{R}$ of our problem, we can use a surrogate model to learn the relationships between the data in the form $y_{i}=f({x}_{i})+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$. The surrogate is commonly a Gaussian Process (GP), a non-parametric model that provides uncertainty quantification 19. GPs are usually defined as $f(x)\sim\mathcal{GP}(0,k({x},{x^{\prime}}))$, placing a zero-mean prior and a covariance given by a kernel function $cov[f({x}),f({x^{\prime}})]=k({x},{x^{\prime}})$ that measures the similarity between inputs. Given a dataset $\mathcal{D}$, the predictive mean and variance are given by the following expressions
@@
where ${k_{x}}=[k({x},{x_{1}}),...,k({x},{x_{n}})]^{\top}\in\mathbb{R}^{n}$, $K\in\mathbb{R}^{n\times n}=(k({x_{i}},{x_{j}}))_{1\leq i,j\leq n}$ and ${y}=[y_{1},...,y_{n}]$. In order to optimize the target problem, after training the surrogate with the available data, a decision is made to select the next points by maximizing the acquisition function. Acquisition functions are heuristics that decide the next point to query by balancing exploration and exploitation. The next inputs are given by
@@
Multi-fidelity BO extends the previous setup by considering an extra input parameter for the surrogate model, the fidelity l. This parameter indicates which source of information the other inputs correspond to (e.g., high fidelity, HF, or low fidelity, LF). In our case, we use a discrete fidelity setting, where $l\in[l_{0},l_{1},...l_{m}]$ and $l_{m}$ corresponds to the highest fidelity among the $m$ levels. In all cases, we limit the analysis to two levels of fidelity ($l_{0}$ and $l_{1}$ and $m=1$) , but the methodology can be extended to more levels. Additionally, these models assume that the cost only depends on the fidelity level $l$ and not on the search space $x$. We employed a modified GP proposed in 10 that extends the previously described surrogate modelling to the multi-fidelity space of $\mathbb{R}^{n+1}$ dimensions, chosen due to its popularity amongst MFBO applications10, 14, 18. The modified kernel is defined as $k(({x},l),({x^{\prime}},l^{\prime}))$, where fidelity is introduced by defining a separate kernel to model the input space, and a kernel to model the correlation and interaction between the different fidelity levels. The mathematical expression is:
@@
Given a dataset of inputs with their associated fidelities and outputs, the model can be trained using standard maximum log-likelihood optimization to obtain the optimal hyperparameters $\lambda_{i}$, $c$ and $\delta$. In the experiments, we set the $l$ values of the HF and LF levels to 1 and 0 respectively following previous studies that used this kernel in a binary setting14, 18.
@@
In the experiments we use two families of acquisition functions, with their single and multi-fidelity versions: Maximum Entropy Search (MES)11 and Expected Improvement (EI)20. We also initially tested Knowledge Gradient (KG)10, but we discarded it due to its computational expense.
In all cases, the experiments are run for 20 independent seeds, and the number of initial samples corresponds to 10% of the total optimization budget, sampled using a Latin hypercube strategy. In the MFBO case, the budget is split to sample 50% of HF and 50% of LF initial points. In the synthetic functions, the total budget corresponds to 50 HF queries. In the chemistry and materials benchmarks, the total budget corresponds to 30 HF queries. Data collection is sequential. All the experiments are run using BOTorch 21 as a standard Bayesian Optimization framework.
@@
Figure 2: Standardized metrics to compare MFBO and SFBO. Simple regret traces are computed using algorithm 1 and provide a direct comparison between the optimization progress in each method by computing the equivalent number of steps from SFBO in the MFBO trace. Discount ($\Delta$) reflects the savings in optimization cost by computing the normalized difference in budget spent by each algorithm ($\texttt{b}^{\rm sf}(\tilde{r}^{*})$ and $\texttt{b}^{\rm mf}(\tilde{r}^{*})$) to find $\tilde{r}^{*}$ (horizontal arrow, see equation 1).
Due to the lack of standardized metrics in the MFBO field17, we define a standardized simple regret and a discount metric to quantify the performance of MFBO over SFBO. Let $f$ be the HF target problem. At a given step $t$ of the SFBO run, the simple regret is defined as $r_{t}=f^{*}-f^{*}_{t}$, where $f^{*}$ is the known global optimum and $f^{*}_{t}$ is the best value of $f$ found so far. The simple regret is computed only at the highest fidelity level in the multi-fidelity runs. To compare multi-fidelity and single-fidelity runs it is necessary to find the corresponding high-fidelity simple regret in the multi-fidelity setting, which can query both the low-fidelity or high-fidelity at a given step. We define a standard algorithm to identify the corresponding HF simple regret at a given budget of an MFBO campaign. The algorithm aligns the ${r}$ values from the MFBO to a common cost scale given by the SFBO cost steps. This makes it possible to compare the performance of the MFBO method at each step of the associated SFBO run. Algorithm 1 explains in detail the procedure to compute simple regret both for single-fidelity and multi-fidelity runs.
@@
We also propose a discount $\Delta$ metric to estimate the savings provided by the MFBO method compared to the SFBO (see figure 2 for a graphical explanation of this metric). The discount variable reflects the difference of budgets required to reach a reference regret value between the multi-fidelity and single-fidelity methods. Specifically, it is computed over two individual SF and MF runs as follows:
$\Delta=\dfrac{\texttt{b}^{\rm sf}(\tilde{r}^{*})-\texttt{b}^{\rm mf}(\tilde{r}^{*})}{\texttt{b}^{\rm sf}(\tilde{r}^{*})},$ (1)
where $\tilde{r}^{*}$ is the corrected best multi-fidelity regret, defined as $\tilde{r}^{*}=r^{\rm mf}_{t}\tau$. The $\tau$ correction accounts for situations where the MFBO method gets to low regrets faster than the SFBO at the beginning of the optimization, but its performance plateaus in the long term as noted in 17. In our experiments, we set $\tau=2$ as a reasonable adjustment, given that $r^{\rm mf}_{t}$ tends to be small. A positive $\Delta$ indicates cost savings with MFBO, meaning it achieves the minimum regret with a lower budget than SFBO. Conversely, a negative $\Delta$ implies that MFBO was more expensive than SFBO. If the SFBO method is unable to reach $\tilde{r}^{*}$, $\Delta$ is capped at 1. Note that equation 1 compares individual runs started at the same seed as displayed in Figure 2, and we report the average $\Delta$ over all the seeds onwards.
@@
We use cost ratio and informativeness metrics to compute the characteristics of the low-fidelity approximation with respect to the high-fidelity. Cost is characterized by the cost ratio $\rho$, which is obtained by dividing the LF cost and the HF cost. Informativeness is characterized by the $R^{2}$ of the LF approximation. To compute this metric, we uniformly sample 100 points for the given problem and extract their associated values at the HF and LF levels. Then, the $R^{2}$ between a linear fitting of the LF prediction and the true HF value is computed. Supplementary Information A.3 shows the plots with the sampled points and the computed $R^{2}$ of each of problems used in this study.
@@
Synthetic functions are simulated black-box problems where a mathematical expression is used to generate points from a surface to be optimized. In the multi-fidelity case, there are lower accuracy approximations of the target function at a lower cost. These can either have a fixed expression or be generated by biasing the original function using a given parameter $\alpha$. In our case, we use the Branin18 function previously used in MFBO works. We also modify a Park function used in MFBO22 by incorporating a parameter $\alpha$ to modulate the bias with a similar behaviour to the Branin case. The expressions of the functions can be found in the Supplementary Information A.1. The cost of querying the high-fidelity level is set to a value of 1, and the cost of the low-fidelity approximations is set to a fraction of this value.
@@
We use or adapt previously reported benchmarks for real experimental optimization problems in the chemistry domain. The Covalent Organic Frameworks (COFs) dataset was used in a previous work on MFBO14, and consists of 608 candidate COFs encoded in a 14-dimensional vector accounting for their composition and crystal structure. The high-fidelity simulation uses a Markov chain Monte Carlo simulation to compute the adsorption of Kr and Xe in the material, and the associated Kr/Xe selectivity. This simulation is expensive, with an average running time of 230 minutes. In this case, a low-fidelity approximation of the selectivity can be obtained using Henry’s law, reducing the computing time to 15 minutes, giving a $\rho$ of 0.065 for this problem.
The second benchmark is derived from the Alexandria library 23, and it was used in previous work on MFBO for materials discovery13. It comprises 1134 molecules encoded using RDKit 2D-descriptors 24 and reduced to a 10-D vector using PCA. The high-fidelity is the experimental polarizability, whereas the low-fidelity is the computed polarizability at the Hartree-Fock 6-31G+ level of theory. $\rho$ is set up to 0.167 in this case following the previous study.
The last benchmark is extracted from the FreeSolv library25. It comprises 641 molecules, encoded using RDKit 2D-descriptors and reduced to a 10-D vector using PCA. The high-fidelity is the experimental free solvation energy, whereas the low-fidelity is the computed solvation energy using molecular dynamics. $\rho$ is set up to 0.1 in this case.
@@ Results
We test the performance of MFBO methods over their SFBO counterparts on two different synthetic functions, Branin and Park (with 2 and 4 dimensions respectively), on two different scenarios (favorable and unfavorable). Synthetic functions are commonly employed as black-box problems to test the performance of BO algorithms (see Synthetic functions). In the case of MFBO, the output of the functions can be biased using a parameter $\alpha$ to provide lower accuracy information sources. This parameter is related to the informativeness of the LF source (in our case, we quantify the informativeness of the LF by computing the $R^{2}$ with respect to the HF, see Metrics). For the favorable scenario, we tune the $\alpha$ parameter to provide a highly informative LF source (LF $R^{2}>0.9$). We set the cost of querying a HF point to 1 and the cost of a LF point to 0.1 ($\rho$ = 0.1). Figure 3 shows the results for the favorable scenario. In this case, both MFBO methods outperform their SFBO counterparts, with maximum discounts of 0.5 and 0.58 for Branin and Park respectively. The MFBO runs get to lower regrets quicker, spending less resources by exploiting the use of LF and HF sources of information. The resulting optimization runs are therefore desirable as they effectively leverage the access to the LF to guide the optimization of the HF level.
@@
Although under the previous settings MFBO provides a higher performance than SFBO, a change in the LF conditions dramatically affects the result of the optimization. In the unfavorable scenario, we set the cost of the LF source to 0.5 (only half of the HF cost, $\rho$ = 0.5) and decrease the informativeness of the functions ($R^{2}$ $<$ 0.75). In this case, the performance of both methods is similar (minimum $\Delta$ of -0.07 and -0.1 for Branin and Park respectively), and the MFBO does not provide a quicker convergence to low regret values. The previous results illustrate how MFBO performance is affected by the cost and the informativeness of the LF source, losing its advantage over SFBO if the LF source is not informative and cheap enough.
@@
Figure 3: MFBO results are dependent on problem conditions. In the favorable scenario (highly informative and cheap LF source, $\rho$ = 0.1, $R^{2}$ > 0.9), MFBO offers better performance than SFBO on the synthetic functions (maximum $\Delta$ of 0.5 and 0.58 for Branin and Park, respectively). The optimization traces show how lower regrets are found at lower budgets under this scenario. When the conditions are changed and a low informative and more expensive LF is used ($\rho$ = 0.5, $R^{2}$ < 0.75), MFBO performance is degraded. Under this scenario, the optimization traces are similar, and MFBO does not offer an advantage over SFBO (minimum $\Delta$ of -0.07 and -0.1 for Branin and Park, respectively).
@@
We translate the previous recommendations to three benchmarks in chemistry and materials science to showcase their utility in real experimental situations. The tasks correspond to scenarios where a cheap and informative LF source is available under the proposed guidelines (see Methods for a detailed explanation of each benchmark). Figure 6a illustrates the results of MFBO and SFBO in these scenarios. In all cases, the MFBO method can get to lower regrets at a lower cost than the SFBO by leveraging the cheaper source of information at hand. In the COFs benchmark (problem settings: $\rho$ = 0.065, $R^{2}$ = 0.98), the MFBO gets to the COF with the highest computed Kr/Xe selectivity in the allocated budget, whereas the SFBO does not find it. In the polarizability benchmark (problem settings: $\rho$ = 0.167, $R^{2}$ = 0.99), the MFBO provides a maximum discount of 0.6 over the SFBO. In this case, the use of a cheap Hartree-Fock computation is successfully combined with the experimental measurements to find the molecule with the highest polarizability. In the solvation energy benchmark (problem settings: $\rho$ = 0.1, $R^{2}$ = 0.88), the MFBO finds the molecule with the highest solvation energy but the SFBO not, as in the COFs case. This example also integrates computed values with experimental measurements to reduce the spent budget. Figure 6b shows the proportion of queries to each fidelity level in the different tasks. By distributing the available budget between the HF and LF levels, the MFBO methods can integrate the cheaper information of the LF approximations to improve the optimization results. The number of calls to each query depends on the specific acquisition function and task. However, in all cases the proportion of calls to the HF level is lower than 0.4, suggesting that this proportion can provide a threshold for a successful MFBO optimization. In these real scenarios, the MFBO methods can exploit the low accuracy but cheap LF levels by querying them more often than the HF level.
@@
Figure 6: Molecules and materials benchmark results. a) Optimization performance for the MFBO and the SFBO methods in each benchmark (see Methods for a description of each task). The maximum $\Delta$ in each case is 1, 1 and 0.67 for the COFs, solvation energy and polarizabiliry tasks, respectively. In all the cases, MFBO is able to achieve lower regrets than SFBO by spending less budget. b) Proportion of queries to each fidelity source in the MFBO setting. The MFBO method balances information gain and budget spent by querying more often the LF values (between 0.6 and 0.8 in the ratio of queries) while also using information from HF values. c) Experimental negative example. Molecule polarizability benchmark conditions were changed so the LF source was expensive and less informative. The resulting $\Delta$ is 0 and 0.03 for the MES and EI families respectively, showing how unfavorable conditions can override the MFBO advantage over SFBO in a real experimental setting.

### v2

@@ Abstract
Although MFBO has garnered interest in the past decade within the ML community, leading to a myriad of model definitions and acquisition policies 9, 10, 11, 12, researchers have only lately started to integrate them within their design methodologies. In the chemical domain, there has been recent interest in incorporating cost awareness into the BO loop13, 14. Regarding the specific multi-fidelity approach, several studies have successfully applied MFBO to the materials discovery domain 15, 16, 17, 18, 19, 20. However, there is a fundamental lack in the assessment of MFBO performance, and due to this, each work represents MFBO performance in different ways 21, 16, 19, 20. Although some metrics have been proposed 15, 19, the lack of clarity and unified criteria on how to assess the benefits of MFBO in the chemical domain has hindered its widespread adoption among the experimental community22. Several factors come into play when assessing its impact, yet the application of MFBO can also be detrimental to the overall optimization process 23. It has been also shown how in the long run the advantage of MFBO over SFBO can be lost 22. Therefore, it is crucial to unify and provide a reliable method to decide when MFBO is better than SFBO.
@@
Figure 1: Multi-fidelity Bayesian Optimization (MFBO) combines expensive but informative, and cheap but approximate sources of information with a ML model to optimize black-box problems at an overall reduced budget. The high-fidelity source (HF, e.g., real lab experiment) provides accurate information but at a high cost. The low-fidelity source (LF, e.g., computer simulation of real experiment) represents a cheaper approximation of the HF. In the illustration, we show the surfaces corresponding to the Branin function used in this study at the HF and one LF level, where the optima are located in different places. The MFBO loop iteratively selects HF or LF to maximize the information gain while reducing the overall optimization cost.
@@ Methods
where f is the target black-box function and $\mathcal{X}$ is the space of all possible input candidates. Considering $\mathcal{D}=[(x_{1},y_{1}),(x_{2},y_{2}),...(x_{n},y_{n})]$ a dataset containing pairs of inputs-outputs $({x}_{n},y_{n})\in\mathbb{R}^{d}\times\mathbb{R}$ of our problem, we can use a surrogate model to learn the relationships between the data in the form $y_{i}=f({x}_{i})+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$. The surrogate is commonly a Gaussian Process (GP), a non-parametric model that provides uncertainty quantification 24. GPs are usually defined as $f(x)\sim\mathcal{GP}(0,k({x},{x^{\prime}}))$, placing a zero-mean prior and a covariance given by a kernel function $cov[f({x}),f({x^{\prime}})]=k({x},{x^{\prime}})$ that measures the similarity between inputs. Given a dataset $\mathcal{D}$, the predictive mean and variance are given by the following expressions
@@
where ${k_{x}}=[k({x},{x_{1}}),...,k({x},{x_{n}})]^{\top}\in\mathbb{R}^{n}$, $K\in\mathbb{R}^{n\times n}=(k({x_{i}},{x_{j}}))_{1\leq l_{0},l_{1}\leq n}$ and ${y}=[y_{1},...,y_{n}]$. In order to optimize the target problem, after training the surrogate with the available data, a decision is made to select the next points by maximizing the acquisition function. Acquisition functions are heuristics that decide the next point to query by balancing exploration and exploitation. The next inputs are given by
@@
Our multi-fidelity BO extends the previous setup by considering an extra input parameter for the surrogate model, the fidelity l. This parameter indicates which source of information the other inputs correspond to (e.g., high fidelity, HF, or low fidelity, LF). In our case, we use a discrete fidelity setting, where $l\in[l_{0},l_{1},...l_{m}]$ and $l_{m}$ corresponds to the highest fidelity among the $m$ levels. In all cases, we limit the analysis to two levels of fidelity ($l_{0}$ and $l_{1}$ and $m=1$) , but the methodology can be extended to more levels. Additionally, these models assume that the cost only depends on the fidelity level $l$ and not on the search space $x$. We employed a modified GP proposed in 11 that extends the previously described surrogate modelling to the multi-fidelity space of $\mathbb{R}^{n+1}$ dimensions, chosen due to its popularity amongst MFBO applications11, 16, 23. The modified kernel is defined as $k(({x},l),({x^{\prime}},l^{\prime}))$, where fidelity is introduced by defining a separate kernel to model the input space, and a kernel to model the correlation and interaction between the different fidelity levels. The mathematical expression is:
@@
Given a dataset of inputs with their associated fidelities and outputs, the model can be trained using standard maximum log-likelihood optimization to obtain the optimal hyperparameters $\lambda_{i}$, $c$ and $\delta$. In the experiments, we set the $l$ values of the HF and LF levels to 1 and 0 respectively following previous studies that used this kernel in a binary setting16, 23.
@@
In the experiments, we use two families of acquisition functions, with their single and multi-fidelity versions: Maximum Entropy Search (MES)12 and Expected Improvement (EI)25. We also initially tested Knowledge Gradient (KG)11, but we discarded it due to its computational expense.
In all cases, the experiments are run for 20 independent seeds, and the number of initial samples corresponds to 10% of the total optimization budget, sampled using a Latin hypercube strategy. In the MFBO case, the budget is split to sample 50% of HF and 50% of LF initial points (this proportion was selected based on preliminary experiments that showed how this ratio did not affect the final MFBO performance). In the synthetic functions, the total budget corresponds to 50 HF queries. In the chemistry and materials benchmarks, the total budget corresponds to 30 HF queries. Data collection is sequential. All the experiments are run using BOTorch 26 as a standard Bayesian Optimization framework.
@@
Recent work in MFBO for materials discovery has seen the emergence of several metrics to evaluate method performance. To measure the progress of the optimization campaign, regret15, 16, Active Learning Metrics (ALM)19 and campaign efficiency20 have been proposed. In the ML literature, regret is the most common metric, although its specific implementation for MFBO cases is not defined21, 11, 23, 18. To measure absolute MFBO performance, that is, how a specific MFBO run compares to its SFBO counterpart under specific experimental conditions, Acceleration Factor (AF) and Enhancement Factor19, cost difference to discover materials in the 99th best percentile15 and MF advantage20 have been used. Importantly, MFBO performance is also dependent on the available budget, and this absolute metric has to somehow capture this dependency. Only some works have explicitly mentioned 20, 16 or studied15 how low-fidelity informativeness affects the final result, quantifying it using the $R^{2}$ or correlation coefficient between the HF and LF source, respectively. Last, some recent works highlighted the need of better metrics to assess MFBO performance 22, 27. To solve this lack of standardization, and building on the previous works, we propose two key metrics - MFBO regret (r) and discount ($\Delta$) to study MFBO performance under different scenarios. We also define $\rho$ and $R^{2}$ as the metrics to characterize the LF source.
@@
Figure 2: Standardized metrics to compare MFBO and SFBO. Simple regret traces are computed using algorithm 1 and provide a direct comparison between the optimization progress in each method by computing the equivalent number of steps from SFBO in the MFBO trace. Discount ($\Delta$) reflects the savings in optimization cost by computing the normalized difference in budget spent by each algorithm ($\texttt{b}^{\rm sf}(\tilde{r}^{*})$ and $\texttt{b}^{\rm mf}(\tilde{r}^{*})$) to find $\tilde{r}^{*}$ (horizontal arrow, see equation 1). The plot is zoomed in to focus on the $\Delta$ score.
We define a standardized simple regret and a discount metric to quantify the performance of MFBO over SFBO along the optimization. This metric aims to inform how the MFBO optimization progresses with respect to a known reference single-fidelity source. This approach is similar to the idea of ALM introduced in 19 (note that it is an "after-the-fact" metric, meaning that the best candidate must be known beforehand, although this is common to all the benchmarks). Let $f$ be the HF target problem. At a given step $t$ of the SFBO run, the simple regret is defined as $r_{t}=f^{*}-f^{*}_{t}$, where $f^{*}$ is the known global optimum and $f^{*}_{t}$ is the best value of $f$ found so far. The simple regret is computed only at the highest fidelity level in the multi-fidelity runs. To compare multi-fidelity and single-fidelity runs it is necessary to find the corresponding high-fidelity simple regret in the multi-fidelity setting, which can query both the low-fidelity or high-fidelity at a given step. We define a standard algorithm to identify the corresponding HF simple regret at a given budget of an MFBO campaign. The algorithm aligns the ${r}$ values from the MFBO to a common cost scale given by the SFBO cost steps. This approach makes it possible to compare the performance of the MFBO method at each step of the associated SFBO run. Algorithm 1 explains in detail the procedure to compute simple regret both for single-fidelity and multi-fidelity runs. Even if in our work we only evaluate two sources of fidelity, the advantage of this regret computation is that it allows to incorporate as many discrete LF sources as the user wants. A potential disadvantage may be the need of a reference optimization case (the SFBO run in this case), but we consider this acceptable as the final objective of the study is to decide when MFBO is better than SFBO.
@@
We also propose a discount $\Delta$ metric to estimate the savings provided by the MFBO method compared to the SFBO (see figure 2 for a graphical explanation of this metric). This metric is inspired is the Acceleration Factor (AF) previously proposed19. $\Delta$ reflects the difference in budgets required to reach a reference regret value between the multi-fidelity and single-fidelity methods. Specifically, it is computed over two individual SF and MF runs as follows:
$\Delta(\tilde{r}^{*})=\dfrac{\texttt{b}^{\rm sf}(\tilde{r}^{*})-\texttt{b}^{\rm mf}(\tilde{r}^{*})}{\texttt{b}^{\rm sf}(\tilde{r}^{*})},\text{where }\tilde{r}^{*}=r^{\rm sf}_{\rm max}-(r^{\rm sf}_{\rm max}-r^{\rm sf}_{\rm min})\tau.$ (1)
$\tilde{r}^{*}$ is the corrected best single-fidelity regret, and $r^{\rm sf}_{\rm max}$ and $r^{\rm sf}_{\rm min}$ are the maximum and minimum regrets in the SFBO run, and $\tau$ is a factor that quantifies the amount of total reduced regret that the user wants to sacrifice in order to get a good MFBO performance. The $\tau$ correction accounts for situations where the MFBO method gets to low regrets faster than the SFBO at the beginning of the optimization, but its performance plateaus in the long term as noted in22. This correction also reflects the dependence of MFBO performance on the available budget, as mentioned in previous work20. We report discounts with $\tau$ = 0.9, although we also investigate the discount obtained for several values of $\tau$ in the synthetic functions benchmark (see SI section A.4. A positive $\Delta$ indicates cost savings with MFBO, meaning it achieves the minimum regret with a lower budget than SFBO. Conversely, a negative $\Delta$ implies that MFBO was more expensive than SFBO. If the MFBO method is unable to reach $\tilde{r}^{*}$, $\Delta$ is set to -1. Note that equation 1 compares individual runs started at the same seed as displayed in Figure 2, and we report the average $\Delta$ over all the seeds onwards.
@@
We use cost ratio and informativeness metrics to compute the characteristics of the low-fidelity approximation with respect to the high-fidelity. Cost is characterized by the cost ratio $\rho$, which is obtained by dividing the LF cost and the HF cost. Informativeness is characterized by the $R^{2}$ of the LF approximation. To compute this metric, we uniformly sample 100 points for the given problem and extract their associated values at the HF and LF levels. Then, the $R^{2}$ between a linear fitting of the LF prediction and the true HF value is computed. Supplementary Information A.5 shows the plots with the sampled points and the computed $R^{2}$ of each of the problems used in this study.
@@
Synthetic functions are simulated black-box problems where a mathematical expression is used to generate points from a surface to be optimized. In the multi-fidelity case, there are lower accuracy approximations of the target function at a lower cost. These can either have a fixed expression or be generated by biasing the original function using a given parameter $\alpha$. In our case, we use the Branin23 function previously used in MFBO works. We also modify a Park function used in MFBO28 by incorporating a parameter $\alpha$ to modulate the bias with a similar behaviour to the Branin case. The expressions of the functions can be found in the Supplementary Information A.2. The cost of querying the high-fidelity level is set to a value of 1, and the cost of the low-fidelity approximations is set to a fraction of this value.
@@
We use or adapt previously reported benchmarks for real experimental optimization problems in the chemistry domain. The Covalent Organic Frameworks (COFs) dataset was used in a previous work on MFBO16, and consists of 608 candidate COFs encoded in a 14-dimensional vector accounting for their composition and crystal structure. The high-fidelity simulation uses a Markov chain Monte Carlo simulation to compute the adsorption of Kr and Xe in the material, and the associated Kr/Xe selectivity. This simulation is expensive, with an average running time of 230 minutes. In this case, a low-fidelity approximation of the selectivity can be obtained using Henry’s law, reducing the computing time to 15 minutes, giving a $\rho$ of 0.065 for this problem.
The second benchmark is extracted from the FreeSolv library29. It comprises 641 molecules, encoded using RDKit 2D-descriptors and reduced to a 10-D vector using PCA. The high-fidelity is the experimental free solvation energy, whereas the low-fidelity is the computed solvation energy using molecular dynamics (MD). $\rho$ is set up to 0.1 in this case, based on an estimation of the difference between running a solvation measurement and an MD simulation.
The last benchmark is derived from the Alexandria library 30, and it was used in previous work on MFBO for materials discovery15. It comprises 1134 molecules encoded using RDKit 2D-descriptors 31 and reduced to a 10-D vector using PCA. The high-fidelity is the experimental polarizability, whereas the low-fidelity is the computed polarizability at the Hartree-Fock 6-31G+ level of theory. $\rho$ is set up to 0.167 in this case following the previous study.
@@ Results
We test the performance of MFBO methods over their SFBO counterparts on two different synthetic functions, Branin and Park (with 2 and 4 dimensions respectively), on two different scenarios (favorable and unfavorable). Synthetic functions are commonly employed as black-box problems to test the performance of BO algorithms (see Synthetic functions). In the case of MFBO, the output of the functions can be biased using a parameter $\alpha$ to provide lower accuracy information sources. This parameter is related to the informativeness of the LF source (in our case, we quantify the informativeness of the LF by computing the $R^{2}$ with respect to the HF, see Metrics). For the favorable scenario, we tune the $\alpha$ parameter to provide a highly informative LF source (LF $R^{2}>0.9$). We set the cost of querying a HF point to 1 and the cost of a LF point to 0.1 ($\rho$ = 0.1). Figure 3 shows the results for the favorable scenario. In this case, both MFBO methods outperform their SFBO counterparts, with maximum discounts of 0.57 and 0.38 for Branin and Park respectively. The MFBO runs get to lower regrets quicker, spending less resources by exploiting the use of LF and HF sources of information. The resulting optimization runs are therefore desirable as they effectively leverage the access to the LF to guide the optimization of the HF level.
@@
Although under the previous settings MFBO provides a higher performance than SFBO, a change in the LF conditions dramatically affects the result of the optimization. In the unfavorable scenario, we set the cost of the LF source to 0.5 (only half of the HF cost, $\rho$ = 0.5) and decrease the informativeness of the functions ($R^{2}$ $<$ 0.75). In this case, the performance of both methods decreases with respect to the favorable scenario. In Branin, the maximum $\Delta$ drops to 0.23, whereas in Park the trend is reversed and the MFBO loses its advantage over SFBO, with a maximum $\Delta$ of -0.44. These results illustrate how MFBO performance is affected by the cost and the informativeness of the LF source, losing its advantage over SFBO if the LF source is not informative and cheap enough. It also reflects how the MFBO setting is problem-dependent and it may be more robust towards problem changing situation
Apart from problem conditions, model choice may also affect MFBO performance. Previous studies have proposed and used many MF models, mainly GPs refs but also DNNs10, 32. We compared the standard BoTorch multi-fidelity model with a MultiTask GP9 to see how model choice can affect the results. The main difference between the two models lies in how fidelities are encoded in the GP. SI section A shows a preliminary study on the fidelity kernel and how both models offer similar performances in regression and MFBO tasks with the Branin function (figures 7 and 8 respectively). Therefore, we stick to the use of the default model provided in BOTorch, focusing on the external factors that can affect the experiments.
@@
Figure 3: MFBO results are dependent on problem conditions. In the favorable scenario (highly informative and cheap LF source, $\rho$ = 0.1, $R^{2}$ > 0.9), MFBO offers better performance than SFBO on the synthetic functions (maximum $\Delta$ of 0.57 and 0.38 for Branin and Park, respectively). The optimization traces show how lower regrets are found at lower budgets under this scenario. When the conditions are changed and a low informative and more expensive LF is used ($\rho$ = 0.5, $R^{2}$ < 0.75), MFBO performance is degraded. Under this scenario, the optimization traces are similar, and MFBO performance is notably reduced (minimum $\Delta$ of 0.23 and -0.44 for Branin and Park, respectively).
@@
Although $\Delta$ is reported at a value of $\tau$ = 0.9 for all the cases, we also investigated the effect of $\tau$ on the final discount. In previous works, the MFBO performance was reported to be lower than SFBO at smaller budgets20, and the authors noted that in the long-run MFBO may also lose its advantage over SFBO22. This trend was observed in both functions and AFs, where low values of $\tau$ and a $\tau$ of 1 provided negative discounts (see figure 9 in the SI). Although the temporal dependence of the MFBO is hard to control for the experimenter, due to the uncertainty of when to stop the sampling, the previous figure shows how there exist "sweet spots" where MFBO can exploit the cheap LF source with the available budget to provide an advantage over SFBO.
@@
We translate the previous recommendations to three benchmarks in chemistry and materials science to showcase their utility in real experimental situations. The tasks correspond to scenarios where a cheap and informative LF source is available under the proposed guidelines (see Methods for a detailed explanation of each benchmark). Figure 6a illustrates the results of MFBO and SFBO in these scenarios. In all cases, the MFBO method can get to lower regrets at a lower cost than the SFBO by leveraging the cheaper source of information at hand. In the COFs benchmark (problem settings: $\rho$ = 0.065, $R^{2}$ = 0.98), the MFBO gets to the COF with the highest computed Kr/Xe selectivity in the allocated budget, whereas the SFBO does not find it. In the polarizability benchmark (problem settings: $\rho$ = 0.167, $R^{2}$ = 0.99), the MFBO provides a maximum discount of 0.6 over the SFBO. In this case, the use of a cheap Hartree-Fock computation is successfully combined with the experimental measurements to find the molecule with the highest polarizability. In the solvation energy benchmark (problem settings: $\rho$ = 0.1, $R^{2}$ = 0.88), the MFBO finds the molecule with the highest solvation energy but the SFBO not, as in the COFs case. This example also integrates computed values with experimental measurements to reduce the spent budget. Figure 6b shows the proportion of queries to each fidelity level in the different tasks. By distributing the available budget between the HF and LF levels, the MFBO methods can integrate the cheaper information of the LF approximations to improve the optimization results. The number of calls to each query depends on the specific acquisition function and task. However, in all cases the proportion of calls to the HF level is lower than 0.4, suggesting that this proportion can provide a threshold for a successful MFBO optimization. A previous study suggested a proportion of LF-HF points of 5:1 as the optimal case for good MFBO performance20, as is the case for the COFs and polarizability benchmarks. In these real scenarios, the MFBO methods can exploit the low accuracy but cheap LF levels by querying them more often than the HF level.
@@
Figure 6: Molecules and materials benchmark results. a) Optimization performance for the MFBO and the SFBO methods in each benchmark (see Methods for a description of each task). The maximum $\Delta$ in each case is 0.68, 0.45 and 0.56 for the COFs, solvation energy and polarizabiliry tasks, respectively. In all the cases, MFBO is able to achieve lower regrets than SFBO by spending less budget. b) Proportion of queries to each fidelity source in the MFBO setting. The MFBO method balances information gain and budget spent by querying more often the LF values (between 0.6 and 0.8 in the ratio of queries) while also using information from HF values. c) Experimental negative example. Molecule polarizability benchmark conditions were changed so the LF source was expensive and less informative. The resulting $\Delta$ is -0.07 and 0.13 for the MES and EI families respectively, showing how unfavorable conditions can override the MFBO advantage over SFBO in a real experimental setting.